a car is moving with high velocity suddenly a wall appears to driver of car, in this case what he should do,


 


i)  go forward and apply brakes as hard as he can


ii)  take sharp turn without touching brakes


iii)  take turn with brakes applied hard


iv)   simply close his eyes and pray to god


                 plz give reason for each statements for why n why not ..   i"ll surely approve

Grade:11

5 Answers

iit jee
13 years ago

hi,

i think he should choose option i),

becoz if he goes according to option ii) and iii) then backside of the car would hit the wall.....................

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Yes definitely the correct option is i)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Sanika Phatak
13 years ago

If the man takes a sharp turn the car will move in a circular path but as it has a high velocity by the formula acc=v^2/R the radius will be large so there is a huge possibility that he will crash.

  Even if he applys brakes and takes a sharp turn it is  the same as above as the acc cant exceed v^2.

If he only applys brakes hard enough, he will stop so option 1 is correct.

 

Please approve if correct!!

Tejas Paresh Lodaya
13 years ago

i) is the correct ans.

1/2mv2=Fx ;x=mv2/2F ;The collision would be avoided if x<_r.

or mv2<_r or F>_mv2/2r. Again ,F=mv2/R or R=mv2/F;

R<_r, F>_MV2/r.

Clearly less force will be involved when brakes are applied.

Plz approve...

dewansh parashar
13 years ago

hi,

i think option 3 is correct.....taking the turn will divert the kinetic energy a bit and applying breaks will avoid his head on collision ,also as frictional force increases on increase in sur.area of body,the frictional force will be more against the velocity if the car moves sideways...

hope my logic is correct....and it helps u...

just a mere guess...

thanku